# Invariance under Structure Translation as the Origin of Host Immune Capacity Conservation from Noether’s Theorem

### v2

@@ 4.1 Spatial Continuous Symmetry
Our starting point is the Lagrangian formulation of the immunological dynamics. The state of the system is described by the generalized coordinates $(q_{1},q_{2},q_{3})$, the system’s dynamics are encoded in a Lagrangian $L=T(\dot{q_{i}})-V(q_{i})-V_{\text{couple}}(q_{1},q_{2})$. During a single immune recognition event, the antigenic structure $q_{2}$ is fixed, and the immune state is held constant due to the conservative approximation; consequently, only the kinetic energy $T$ and the coupling potential $V_{\text{couple}}$ evolve over time. The actual path taken by the system between two times, and is the one that minimizes the action $S=\int L\,\mathrm{d}t$.
The symmetry can be explained by an invariance in the system’s dynamics. We therefore postulate that the action functional $S$ remains invariant under a continuous transformation involving the coupling between antigenic epitope structure and the corresponding immune repertoire. Specifically, after imposing the coupling constraint $q1=kq_{2}$, we consider a transformation that smoothly translates the system in the $q_{2}$ (and $q_{1}$) coordinate by an infinitesimal amount $\epsilon:q_{1}\to q_{1}+k\epsilon,q_{2}\to q_{2}+\epsilon$, The symmetry principle is mathematically expressed by the requirement that the Lagrangian remains invariant under this transformation:
$L=L(q_{1}+k\epsilon,q_{2}+\epsilon,q_{3},\dot{q_{1}},\dot{q_{2}},\dot{q_{3}})$ (6)
This is not an a priori assumption, but the simplest dynamical explanation for the observed phenomenological degeneracy.
@@
$\delta S=\int_{t_{1}}^{t_{2}}\left(\frac{\partial L}{\partial q_{1}}(k\epsilon)+\frac{\partial L}{\partial q_{2}}(\epsilon)+\frac{\partial L}{\partial\dot{q}_{1}}\cdot 0+\frac{\partial L}{\partial\dot{q}_{2}}\cdot 0\right)\mathrm{d}t=0$ (8)
$\delta S=\epsilon\int_{t_{1}}^{t_{2}}\left(k\frac{\partial L}{\partial q_{1}}+\frac{\partial L}{\partial q_{2}}\right)\mathrm{d}t=0$ (9)
Since $\delta S=0$, it follows that
$k\frac{\partial L}{\partial q_{1}}+\frac{\partial L}{\partial q_{2}}=\frac{\mathrm{d}}{\mathrm{d}t}(kp_{1}+p_{2})=0$ (10)
The equation implies that the quantity $I_{\mathrm{Spatial}}\equiv k*p_{1}+p_{2}$ is conserved.
After the short immune recognition progress, the generalized force $Q$ begins to affect the system, its introduction necessarily breaks the strict conservation of $I_{\mathrm{Spatial}}$, as in equation (11). While $Q$ supplies a non-conservative impulse that changes the value of $I_{\mathrm{Spatial}}$, its functional form $Q(q_{1})$ is itself invariant under the symmetry transformation. Consequently, the symmetry-induced equivalence between the two antigenic states is preserved throughout the dynamical evolution, even in the absence of a conserved quantity. The state of the system at $t_{2}$ is characterized by the Lagrangian in Equation (5), and the change of $I_{\mathrm{Spatial}}$ is:
$\frac{\mathrm{d}I_{\mathrm{Spatial}}}{\mathrm{d}t}=k\frac{\mathrm{d}p_{1}}{\mathrm{d}t}+\frac{\mathrm{d}p_{2}}{\mathrm{d}t}=k\frac{\partial L}{\partial q_{1}}+\frac{\partial L}{\partial q_{2}}+kQ(q_{1})=kQ(q_{1})$ (11)
$\Delta I_{\mathrm{Spatial}}=\int_{t_{1}}^{t_{2}}\frac{\mathrm{d}I_{\mathrm{Spatial}}}{\mathrm{d}t}\mathrm{d}t=\int_{t_{1}}^{t_{2}}kQ(q_{1})\mathrm{d}t=k\int_{t_{1}}^{t_{2}}Q(q_{1})\mathrm{d}t$ (12)
$I_{\mathrm{Spatial}}(t_{2})=kp_{1}(t_{1})+p_{2}(t_{1})+k\int_{t_{1}}^{t_{2}}Q(q_{1})\mathrm{d}t=I_{\mathrm{Spatial}}(t_{1})+k\int_{t_{1}}^{t_{2}}Q(q_{1})\mathrm{d}t$ (13)
@@ 4.2 Temporal Continuous Symmetry
Through the application of analytical mechanics, we have derived a conserved quantity, $I_{\mathrm{Spatial}}=kp_{1}(t_{1})+p_{2}(t_{1})+k\int_{t_{1}}^{t_{2}}Q(q_{1})\mathrm{d}t$, and an invariant $I_{\mathrm{Temporal}}\equiv\int_{t_{1}}^{t_{2}}Q(q_{1})\dot{q}_{i}\ \mathrm{d}t$ directly from the symmetry inherent in the immune recognition process. This mathematical result provides a fundamental constraint governing the dynamics of the host-pathogen system. The existence of these conservation law is independent of specific pathological contexts; it is a general principle arising from the postulated structure of the interaction Lagrangian. The critical task that follows is to elucidate the profound biological implications of this constraint for immunology.
5 The Static Properties and Dynamic Responses of the Immune Capacity
@@
$I_{\mathrm{Spatial}}=kp_{1}(t_{1})+p_{2}(t_{1})+k\int_{t_{1}}^{t_{2}}Q(q_{1})\mathrm{d}t$ (21)
$I_{\mathrm{Temporal}}\equiv\int_{t_{1}}^{t_{2}}Q(q_{1})\dot{q}_{i}\ \mathrm{d}t$ (22)
@@ 5.1 The Dimensionality of Immune Capacity
To determine the dimensions of the Immune Capacity $\mathbf{I}$, we must first define the dimensions of the generalized coordinate $q_{1}$ and $q_{2}$. We posit that $q_{1}$ shows the diversity of immune repertoire and is dimensionless; $q_{2}$ is also an emergent, dimensionless order parameterthat integrates key physicochemical properties of the antigenic epitope, such as charge distribution, hydrophobicity patterns, and steric topography, into a single quantitative measure of ntigenic identity. The coupling constant $k$ is also dimensionless. Consequently, a translation in antigenic space, $q_{1}\to q_{1}+k\epsilon$, $q_{2}\to q_{2}+\epsilon$, is a dimensionless operation.
@@
$I_{\mathrm{Spatial}}]=[p]=\frac{[E]\times[t]}{[q_{2}]}=\frac{[\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-2}\times[\mathrm{T}]}{1}=[\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-1}$ (23)
The dimension of the generalized force $Q$ is dictated by the principle of virtual work, $\delta W=Q\delta q1$ must possess the dimensions of energy ($\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-2}$), Consequently, the dimensions of $Q$ are constrained by those of the corresponding virtual displacement $\delta q1$, Specifically, if the generalized coordinate $q_{1}$ is dimensionless, the associated generalized force $Q$ must itself carry the dimensions of energy to ensure dimensional consistency of the virtual work expression. It therefore follows that $I_{\mathrm{Temporal}}$ possesses the dimensions of energy.
@@ 5.2 The Role of Innate Immunity within the Conservation
Within the established analytical mechanics framework of the immune response, characterized by the quantity $\mathbf{I}(I_{\mathrm{Spatial}},I_{\mathrm{Temporal}})$, innate immunity and other host systems are not represented by explicit dynamical variables. Instead, their influence is modeled as a modulation of key system parameters. Specifically, a robust innate immune response modulates the coupling constant $k$, the antigenic momentum $p_{2}$ and generalized force $Q(q_{1})$. It first diminishes the effective antigenic load, thereby reducing the pathogenic impulse $p_{2}$ that the adaptive immune system must counter, then it enhances the efficiency of antigen presentation and co-stimulatory signaling, which is quantified by an increase in the coupling constant $k$. Concurrently, host metabolic and signaling activities underpin the generalized force $Q(q_{1})$, which represents the energetic cost required to maintain immune homeostasis and the degree of deviation from this steady state during disease. Thus, the dynamical evolution governed by the conservation law $\mathbf{I}(I_{\mathrm{Spatial}},I_{\mathrm{Temporal}})$ is fundamentally shaped by the parameters $k$, $p_{2}$ and $Q$, These parameters collectively establish the initial conditions and determine the operational trajectory of the immune response along the constraint defined by the conservation law.
5.3 Static Analysis of Homeostatic State
Under homeostatic conditions in a healthy individual, the immune system maintains a dynamic equilibrium characterized by continuous low-level exposure to environmental antigens. In this steady state, the generalized forces $Q$ is approximately zero, indicating a balance between energy dissipation and input. Here, the conserved quantity is expressed as $I_{\mathrm{Spatial}}=kp_{1}+p_{2}$, while $I_{\mathrm{Temporal}}=0$, reflecting that the temporal variation arises primarily from the baseline activity of the adaptive immune system. The antigenic momentum $p_{2}$ remains constant at a low magnitude, with $\Delta p_{2}=0$. This baseline activity defines the system’s intrinsic immune capacity. A robust, young system exhibits a high value of $k$ and $m_{1}$ and resulting in a low $\dot{(}q_{1})$, which signifies strong dynamic stability. In contrast, an aged or immunosenescent system shows reduced baseline activity, manifesting as decreased $k$ and $m_{1}$ and an elevated total $\dot{(}q_{1})$. Since $p_{2}$ is fixed under steady-state conditions, the value of $I_{\mathrm{Spatial}}$ is proportional to the magnitudes of $k$ and $m_{1}$.
$I_{\mathrm{Spatial}}\propto km_{1}$ (25)
@@ 5.4 The Immune Dynamics under Impulse-Response Perturbation
The introduction of a significant pathogenic challenge disrupts this equilibrium, imposing a positive impulse $\Delta p_{2}>0$. $I_{\mathrm{Spatial}}$ mandates an instantaneous compensatory adjustment such that $\Delta I_{\mathrm{Spatial}}=k\Delta p_{1}+\Delta p_{2}=0$, leading to the fundamental constraint:
$k\Delta p_{1}=-\Delta p_{2}$ (26)
Theoretically, this constraint predicts a negative generalized velocity $\dot{q}_{1}$ and a consequent reduction in $q_{1}$. This prediction is consistent with published evidence showing that immune diversity during disease states is lower than that in healthy conditions.
This equation formalizes the zero-sum nature of the host-pathogen interaction. The magnitude of the immune system’s observed response, quantified by the change in the generalized velocity $\Delta\dot{q_{1}}$, is therefore determined by the magnitude of the pathogenic impulse $\Delta p_{2}$and the system’s pre-existing total $I_{\mathrm{Spatial}}$. A large $\Delta p_{2}$ necessarily forces a substantial change in the immune state.
The host can achieve the requisite compensatory change $k\Delta p_{1}$ through several mechanistic strategies that modulate the parameters of the conservation equation. These include modulating the coupling efficiency $k$ through enhanced antigen presentation or co-stimulation, altering the effective mass $m_{1}$ via clonal expansion or recruitment, or directly reducing the antigenic load $p_{2}$ to diminish the effective $\Delta p_{2}$.
A robust, healthy system is characterized by a high intrinsic $I_{\mathrm{Spatial}}$, $k$ and $m_{1}$. For a given pathogenic impulse $\Delta p_{2}$, a high $I_{\mathrm{Spatial}}$, $k$ and $m_{1}$ results in a smaller requisite change in the repertoire velocity $\Delta\dot{q_{1}}$; the system can absorb the perturbation with minimal reconfiguration. Conversely, a system with a low $I_{\mathrm{Spatial}}$, $k$ or $m_{1}$ has minimal spare capacity and must undergo a large, potentially disruptive change in $\dot{q_{1}}$ to maintain the conservation law. Consequently, during the transient phase of a response where parameters $k$, $m_{1}$, and $I_{\mathrm{Spatial}}$ can be considered approximately constant, the observed repertoire change rate $\dot{q_{1}}$ is inversely proportional to $I_{\mathrm{Spatial}}$.
The change in momentum is:
$\Delta p_{1}=m_{1}\Delta\dot{q_{1}}=-\Delta p_{2}$ (27)
Therefore, the magnitude of the system’s response is given by:
$\left|\Delta\dot{q_{1}}\right|=\frac{\left|\Delta p_{2}\right|}{km_{1}}\propto\frac{1}{I_{\mathrm{Spatial}}}$ (28)
@@
Based on the antigenic impulse $p_{2}$, the system is driven from its homeostatic equilibrium. This deviation subsequently modulates the generalized force $Q$, which performs work on the immune system, thereby amplifying the positive displacement from equilibrium. The magnitude of this deviation at time $t$ is quantified by $I_{\mathrm{Temporal}}$. As antigen clearance progresses, the antigenic momentum $p_{2}$ declines. Owing to delayed physiological responses such as those mediated by the endocrine system, the generalized force $Q$ can be considered constant over short intervals, while $I_{\mathrm{Spatial}}$ remains conserved. The reduction in $p_{2}$ induces a corresponding increase in $p1$, leading to an increasion of immune diversity and a gradual attenuation of $Q$. Eventually, $Q$ may reverse direction, initiating a restorative phase that diminishes the system’s deviation and ultimately returns it to the baseline homeostatic state.

### v3

@@ 4.1 Spatial Continuous Symmetry
Our starting point is the Lagrangian formulation of the immunological dynamics. The state of the system is described by the generalized coordinates $(q_{1},q_{2},q_{3})$, the system’s dynamics are encoded in a Lagrangian $L=T(\dot{q_{i}})-V(q_{i})-V_{\text{couple}}(G(q_{1},q_{2}))$. During a single immune recognition event, the antigenic structure $q_{2}$ is fixed, and the immune state is held constant due to the conservative approximation; consequently, only the kinetic energy $T$ and the coupling potential $V_{\text{couple}}$ evolve over time. The actual path taken by the system between two times, and is the one that minimizes the action $S=\int L\,\mathrm{d}t$.
The symmetry can be explained by an invariance in the system’s dynamics. We postulate that the action $S$ remains invariant under a continuous transformation involving the coupling between antigenic epitope structure and the corresponding immune repertoire.
Consider a transformation that smoothly translates the system in the $q_{1}$ and $q_{2}$ coordinate by an infinitesimal amount $q_{1}\to q_{1}+\delta q_{1},q_{2}\to q_{2}+\delta q_{2}$, The symmetry principle is mathematically expressed by the requirement that the action remains invariant under this transformation:
$L^{\prime}=L(q_{1}+\delta q_{1},q_{2}+\delta q_{2},q_{3},\dot{q_{1}},\dot{q_{2}},\dot{q_{3}})$ (8)
The requirement of indistinguishability of the system under the transformation implies that the coupling between $q_{1}$ and $q_{2}$ remains unchanged, the value of $G$ is preserved. We therefore introduce such transformation where $\epsilon$ is an infinitesimal amount:
$G(q_{1},q_{2})=G(q_{1}+\delta q_{1},q_{2}+\delta q_{2})$ (9)
$\nabla G\cdot\delta\mathbf{q}=\frac{\partial G}{\partial q_{1}}\delta q_{1}+\frac{\partial G}{\partial q_{2}}\delta q_{2}=0$ (10)
$\delta q_{1}=\epsilon\frac{\partial G}{\partial q_{2}}=\epsilon K_{2},\delta q_{2}=-\epsilon\frac{\partial G}{\partial q_{1}}=-\epsilon K_{1}$ (11)
The variation of the action is
@@
Assuming the transformation is a symmetry, $\delta S=0$, it follows that
$\delta S=\int_{t_{1}}^{t_{2}}\left(\frac{\partial L}{\partial q_{1}}\delta q_{1}+\frac{\partial L}{\partial q_{2}}\delta q_{2}+\frac{\partial L}{\partial\dot{q}_{1}}\cdot 0+\frac{\partial L}{\partial\dot{q}_{2}}\cdot 0\right)\mathrm{d}t=0$ (13)
$\delta S=\epsilon\int_{t_{1}}^{t_{2}}\left(\frac{\partial L}{\partial q_{1}}\frac{\partial G}{\partial q_{2}}-\frac{\partial L}{\partial q_{2}}\frac{\partial G}{\partial q_{1}}\right)\mathrm{d}t=0$ (14)
this condition leads to the conservation law:
$\frac{\partial L}{\partial q_{1}}\frac{\partial G}{\partial q_{2}}-\frac{\partial L}{\partial q_{2}}\frac{\partial G}{\partial q_{1}}=p_{1}K_{2}-p_{2}K_{1}=\text{Constant}$ (15)
$K_{1}K_{2}\left(\frac{p_{1}}{K_{1}}-\frac{p_{2}}{K_{2}}\right)=\text{Constant}$ (16)
where $p_{1}$ and $p_{2}$ are the generalized momenta. This conserved quantity arises from Noether’s theorem applied to the symmetry generated by the coupling function $G$. The equation implies that the quantity $I_{\mathrm{Spatial}}\equiv p_{1}K_{2}-p_{2}K_{1}$ is conserved. This is not an a priori assumption, but the simplest dynamical explanation for the observed phenomenological degeneracy.
After the short immune recognition progress, the generalized force $Q$ begins to affect the system, its introduction breaks the strict conservation of $I_{\mathrm{Spatial}}$.
@@ 4.2 Temporal Continuous Symmetry
Through the application of analytical mechanics, we have derived a conserved quantity, $I_{\mathrm{Spatial}}=K_{2}p_{1}-K_{1}p_{2}$, and an invariant $I_{\mathrm{Temporal}}\equiv\int_{t_{1}}^{t_{2}}Q(q_{1})\dot{q}_{i}\ \mathrm{d}t$ directly from the symmetry inherent in the immune recognition process. This mathematical result provides a fundamental constraint governing the dynamics of the host-pathogen system. The existence of these conservation law is independent of specific pathological contexts; it is a general principle arising from the postulated structure of the interaction Lagrangian. The critical task that follows is to elucidate the profound biological implications of this constraint for immunology.
5 The Properties and Dynamic Responses of the Immune Capacity
@@
$I_{\mathrm{Spatial}}=K_{2}p_{1}-K_{1}p_{2}$ (24)
$I_{\mathrm{Temporal}}=\int_{t_{1}}^{t_{2}}Q(q_{1})\dot{q}_{i}\ \mathrm{d}t$ (25)
@@ 5.1 The Dimensionality of Immune Capacity
To determine the dimensions of the Immune Capacity $\mathbf{I}$, we must first define the dimensions of the generalized coordinate $q_{1}$ and $q_{2}$. We posit that $q_{1}$ shows the diversity of immune repertoire and is dimensionless; $q_{2}$ is also an emergent, dimensionless order parameterthat integrates key physicochemical properties of the antigenic epitope, such as charge distribution, hydrophobicity patterns, and steric topography, into a single quantitative measure of ntigenic identity. Consequently, a translation in antigenic space, $q_{1}\to q_{1}+\delta q_{1}$, $q_{2}\to q_{2}+\delta q_{2}$, is a dimensionless operation.
@@
$I_{\mathrm{Spatial}}]=[p]=\frac{[E]\times[t]}{[q_{1}]}=\frac{[\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-2}\times[\mathrm{T}]}{1}=[\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-1}$ (26)
The dimension of the generalized force $Q$ is dictated by the principle of virtual work, $\delta W=Q\delta q_{1}$ must possess the dimensions of energy ($\mathrm{M}][\mathrm{L}]^{2}[\mathrm{T}]^{-2}$), Consequently, the dimensions of $Q$ are constrained by those of the corresponding virtual displacement $\delta q_{1}$, Specifically, if the generalized coordinate $q_{1}$ is dimensionless, the associated generalized force $Q$ must itself carry the dimensions of energy to ensure dimensional consistency of the virtual work expression. It therefore follows that $I_{\mathrm{Temporal}}$ possesses the dimensions of energy.
@@ 5.2 The Role of Innate Immunity within the Conservation
Within the established analytical mechanics framework of the immune response, characterized by the quantity $\mathbf{I}(I_{\mathrm{Spatial}},I_{\mathrm{Temporal}})$, innate immunity and other host systems are not represented by explicit dynamical variables. Instead, their influence is modeled as a modulation of key system parameters. Specifically, a robust innate immune response modulates the antigen stimuli $q_{2}$, the coupling function $G$ and generalized force $Q(q_{1})$. It first diminishes the effective antigenic load, thereby reducing the pathogenic impact that the adaptive immune system must counter, then it enhances the efficiency of antigen presentation and co-stimulatory signaling, which is quantified by an increase in the coupling function $G$. Concurrently, host metabolic and signaling activities underpin the generalized force $Q(q_{1})$, which represents the energetic cost required to maintain immune homeostasis and the degree of deviation from this steady state during disease. Thus, the dynamical evolution governed by the conservation law $\mathbf{I}(I_{\mathrm{Spatial}},I_{\mathrm{Temporal}})$ is fundamentally shaped by $G$, and $Q$, These parameters collectively establish the initial conditions and determine the operational trajectory of the immune response along the constraint defined by the conservation law.
5.3 Static Properties and Biological Interpretation of the Conserved Quantities
The quantity $I_{\mathrm{Spatial}}$ arises from the translational invariance of the generalized coordinates. It therefore constitutes a conservation law analogous to momentum conservation, quantifying the resistance to altering the immune system’s dynamical state during antigen interaction. An increase in the cell activation barrier $\mu$ or a decrease in immune-network integrity $\eta$ elevates this conserved quantity, reflecting greater difficulty in changing the system’s state. This manifests biologically as impaired immune activation and a sluggish response. Appropriate antigenic stimulation can reduce the magnitude of this conserved quantity, thereby lowering the activation threshold of the immune system, while excessive stimulation drives it too low. An excessively small value of $I_{\mathrm{Spatial}}$ indicates that the system’s state can be altered too readily, often correlating with hypersensitivity or autoimmune reactions. These dynamic adjustments are consistent with existing experimental observations. An optimal, intermediate magnitude representing a balanced level of resistance is therefore associated with appropriate immune responsiveness.
Similarly, $I_{\mathrm{Temporal}}$ originates from temporal translational invariance, leading to a consistency in energy variation. It characterizes the invariance of energy transfer during the immune system’s evolution under a given generalized force $Q$. When this quantity is too low, the open immune system receives an insufficient net energy influx, resulting in a weak response that may fail to clear antigens efficiently and can lead to chronic infection. If the value is too high, the immune reaction becomes excessively vigorous, potentially triggering pathological outcomes such as cytokine storms.
@@ 5.4 The Immune Dynamics under Impulse-Response Perturbation
The introduction of a significant pathogenic challenge imposes a positive impulse $\Delta p_{2}>0$ to the system. $I_{\mathrm{Spatial}}$ mandates an instantaneous compensatory adjustment such that $\Delta I_{\mathrm{Spatial}}=K_{1}\Delta p_{2}-K_{2}\Delta p_{1}=0$, leading to the fundamental constraint:
$K_{1}\Delta p_{2}=K_{2}\Delta p_{1}=\frac{K_{2}N\mu\Delta\dot{q}_{1}}{\eta}$ (28)
$\dot{q}_{1}=\frac{K_{1}\eta}{K_{2}N\mu}\Delta p_{2}$ (29)
Theoretically, this constraint predicts a positive generalized velocity $\dot{q}_{1}$. This prediction is consistent with published evidence showing that activated immune system during severe disease states has a higher expansion rate than that in milder diseases and healthy conditions.
The magnitude of the immune system’s response, quantified by the generalized velocity $\dot{q_{1}}$, is therefore determined by the magnitude of the pathogenic impulse $\Delta p_{2}$ and the system’s pre-existing total $I_{\mathrm{Spatial}}$. A large $\Delta p_{2}$ necessarily forces a substantial change in the immune state.
The host can achieve the requisite compensatory change $k\Delta p_{1}$ through several mechanistic strategies that modulate the parameters of the conservation equation. These include modulating the coupling efficiency $K_{1}$ through enhanced antigen presentation , altering the co-stimulation and inter-cell communication efficiency $\eta$, or directly reducing the antigenic load $p_{2}$ to diminish the effective $\Delta p_{2}$.
Based on the dynamical framework derived above, the change in the immune system’s state velocity $\dot{q}_{1}$, following an antigenic impulse $\Delta p_{2}$, is given by equation (29). Thus, the magnitude of $\dot{q}_{1}$ is directly proportional to the receptor marginal affinity K2, the antigenic impulse $\Delta p1$, and the network structural integrity $\eta$, while inversely proportional to the antigen marginal immunogenicity K1, the total lymphocyte number N, and the single-cell inertial parameter $\mu$. Among these factors, the parameters K1, K2, $\eta$, and $\mu$ typically exhibit relatively small inter-individual variation in a given population, whereas the total lymphocyte count Ncan vary substantially between individuals (e.g., young vs. elderly, healthy vs. immunocompromised). Consequently, a robust, healthy immune system with a large N experiences only a small change in its state velocity for a given pathogenic impulse; such a system can absorb the disturbance with minimal reconfiguration, maintaining homeostasis. In contrast, a weaker system with a smaller Nmust undergo a larger change in its velocity and potentially disruptive (i.e., a more drastic alteration of its repertoire composition or activation state) to satisfy the conservation law. This result predicts that the observed lineage-turnover rate or the speed of immune-repertoire reshaping should be inversely correlated with the overall immune competence, a prediction consistent with clinical observations that immunosenescence or immunodeficiency is often accompanied by exaggerated and dysregulated immune responses to infections.
@@
Based on the antigenic impulse $p_{2}$, the system is driven from its homeostatic equilibrium. This deviation subsequently modulates the generalized force $Q$, which performs work on the immune system, thereby amplifying the positive displacement from equilibrium. The magnitude of this deviation at time $t$ is quantified by $I_{\mathrm{Temporal}}$. As antigen clearance progresses, the antigenic momentum $p_{2}$ declines. The reduction in $p_{2}$ induces a corresponding change of $Q$. Eventually, $Q$ may reverse direction, initiating a restorative phase that diminishes the system’s deviation and ultimately returns it to the baseline homeostatic state.